OpenStudy (anonymous):

Find the slope and the y-intercept of the equation y − 3(x − 1) = 0.

OpenStudy (anonymous):

y-3x+3=0 y-3x=-3 y=3x-3 formula is y=mx+b m being x b being y whats x-intercept and y-intercept?

OpenStudy (anonymous):

m being slop* what is slope*

OpenStudy (anonymous):

y − 2x = 3 -2y + 3x = 1 y + 3x = 2 y + 2x = 3 are the answers